A body traverses 128 m in first four second and 72m in the next six second. If the acceleration remains uniform,then what will be the distance the body traversed after next two second?

Grade:11

1 Answers

Hamza
6 years ago
Your question should read as:What is the acceleration of a body moving 16m in 4s and moving 9m in 3s with uniform acceleration?We shall take rectilinear motion with the equation of motion as follows.x = ut +1/2 at^2,where a is uniform acceleration and u is initial (t=0s) velocity.In case 1 x=16m and t=4s. Therefore, 16=4u+1/2 a(4)^2=4u+8a or 4=u+2a…………….(1)For second case,quite similarly, 3=u+1.5a………………………………………………………..(2)Eq(1)-Eq(2) gives a=2m/s^2.